How can you assume that Sij = 0 for all i and j. For example S12, how can the overlap integral of two adjacent Pz orbitals (one on the carbon and one on the adjacent oxygen) equal zero? Similarly for the other overlap integrals S13=S14=S21=S31=S41 .


1
Expert's answer
2021-04-13T08:03:39-0400

If the wave functions do not overlap, then the overlap integral is zero. The integral can also be zero if the wave functions have positive and negative aspects that cancel out. If the overlap integral is zero, then the wave functions are called orthogonal. As the distance of the functions get closer to R = 0, then the overlap integral gets closer to the maximum overlap integral of S = 1.

The wave functions of atomic orbitals decrease exponentially with distance. Orbital overlap is non-zero when two atoms are close together.

Overall, the full equation for the overlap integral, which is hard to calculate but is explained here, is

\[S(R)= \left \langle 1s_A | 1s_B \right \rangle = e^{-R/a_0} \left (1 + \dfrac{R}{a_0} + \dfrac{R^2}{3a_0^2} \right)\]